OpenStudy (anonymous):

For the functions below, state the vertical translation? y = -sin(x – (π / 4)) + 2

OpenStudy (jdoe0001):

see it now?

OpenStudy (anonymous):

Would it be D = 1.0

OpenStudy (jdoe0001):

1.0?

OpenStudy (jdoe0001):

ahemm one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is my graph

OpenStudy (jdoe0001):

anyway, with a trig expression like sin(WHATEVER) + SOMEVALUE, the vertical shift is at SOMEVALUE, if positive upward, if negative downward

OpenStudy (anonymous):

So if it goes up it will be positive 3.0?

OpenStudy (jdoe0001):

hmmm, let's see, what would be SOMEVALUE from \(\large y=-sin\pmatrix{x-\frac{\pi}{4}}\color{blue}{+2} \ \ ? \)

OpenStudy (jdoe0001):

3.0?

OpenStudy (anonymous):

2? ._.

OpenStudy (anonymous):

I am really confused..

OpenStudy (jdoe0001):

from THAT trig standard form, THAT part of it, is the vertical shift/translation/transformation

OpenStudy (anonymous):

ahhhhh, ok. I think I get it now.

OpenStudy (anonymous):

So for example y = 2cos(2πx) will have 0 for vertical translation, right?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

because it'd be the same as \(\large 2cos(2\pi x)+0\)

OpenStudy (anonymous):

Ohhhhh, ok. I get it! Thank you!! c:

OpenStudy (jdoe0001):

yw